# Learning to Reason Across Parallel Samples for LLM Reasoning

### v1

@@ Appendix B Prompt Details
{mdframed}
SFT Data Construction Prompt: Here is a question and some proposed answers. You need to evaluate each answers individually, check whether each answer directly addresses the original question, assess the correctness of each answer based on logical reasoning, calculations, and accuracy relative to the question. After thorough evaluation, identify one correct answer. If the correct answer is not in the provided proposed answers, the Assistant will combine the correct partial responses to proposed answers and provide the correct answer. Make the reasoning process concise and to the point. The reasoning process and answer are enclosed within <think></think> and <answer></answer> tags, respectively,i.e., <think>reasoning process here</think> <answer>answer here</answer>. Here is the Question and proposed answers: Question and 5 Answers.
Here is the labeled answer with answer tags ’<answer>Ground Truth Answer</answer>’, and most likely the correct answer. If none of the proposed answers is correct, come up with the reasoning process to arrive at the actual correct final answer. **Attention: Do not ever mention the answer is given in the response, but you need to justify it from the problem and proposed answers!**

### v2

@@ Appendix B Prompt Details
SFT Data Construction Prompt: Here is a question and some proposed answers. You need to evaluate each answers individually, check whether each answer directly addresses the original question, assess the correctness of each answer based on logical reasoning, calculations, and accuracy relative to the question. After thorough evaluation, identify one correct answer. If the correct answer is not in the provided proposed answers, the Assistant will combine the correct partial responses to proposed answers and provide the correct answer. Make the reasoning process concise and to the point. The reasoning process and answer are enclosed within <think></think> and <answer></answer> tags, respectively,i.e., <think>reasoning process here</think> <answer>answer here</answer>. Here is the Question and proposed answers: Question and 5 Answers. Here is the labeled answer with answer tags ’<answer>Ground Truth Answer</answer>’, and most likely the correct answer. If none of the proposed answers is correct, come up with the reasoning process to arrive at the actual correct final answer. **Attention: Do not ever mention the answer is given in the response, but you need to justify it from the problem and proposed answers!**